OpenStudy (anonymous):

Christine’s Butter Cookies sells large and small tins of cookies. The factory can prepare at most 200 tins of cookies each day. Each large tin requires 2 pounds of butter. Each small tin requires one pound of butter. Only 300 lbs. of butter is available daily. The profit from each day’s cookie production is estimated by the function, f(x,y) = $6.00x + $4.80y, where x is the number of large tins sold and y is the number of small tins sold. What is the maximum profit that Christine’s Butter Cookies can expect in a day?

OpenStudy (kc_kennylau):

We also have 2x+y=300 since "each large tin requires 2 pounds ...."

OpenStudy (kc_kennylau):

y=300-2x

OpenStudy (whpalmer4):

and \[x+y \le 200\] because of "the factory can prepare at most 200 tins of cookies each day"

OpenStudy (anonymous):

possible answers: $920 $1080 $600 $480

OpenStudy (kc_kennylau):

f(x,y)=6x+4.8y=6x+4.8(300-2x)=1440-3.6x

OpenStudy (kc_kennylau):

So when x=0 the profit would be the greatest

OpenStudy (kc_kennylau):

But x cannot be 0 or otherwise y would be 300

OpenStudy (kc_kennylau):

So we minimize the x

OpenStudy (kc_kennylau):

And you'd get x=100

OpenStudy (kc_kennylau):

Plug it back into 1440-3.6x

OpenStudy (kc_kennylau):

And you should get the answer

OpenStudy (anonymous):

that isn't right, it would be 80

OpenStudy (whpalmer4):

What would be 80?

OpenStudy (kc_kennylau):

1440-3.6*100 doesn't seem to be 80?

OpenStudy (anonymous):

oh just kidding. i didn't type it correctly. it is 1080. that is right then.

OpenStudy (kc_kennylau):

Do you understand why x=100?

OpenStudy (kc_kennylau):

2x+y=300 x+y=200

OpenStudy (anonymous):

oh, okay. I get it. Thank you @kc_kennylau

OpenStudy (kc_kennylau):

no problem